OpenStudy (anonymous):

Hello OpenStudy! What is a power of 10? Be able to identify 3 numbers that can be written as a power of ten or is written as a power of 10. Be sure that one example is a fraction. If anyone can help me on this, thanks~!

OpenStudy (anonymous):

@Ashleyisakitty @anteater @AravindG Anyone of you can help me, thank you!

OpenStudy (anonymous):

power of 10: \[10^x\] for any x \[10^0, 10^1, 10^2, 10^3,10^4, ..., etc = 1, 10, 100, 1000, 10000,...,etc\]

OpenStudy (anonymous):

also true for: \[10^{-1}, 10^{-2}, 10^{-3} = \frac{ 1 }{ 10 }, \frac{ 1 }{ 100 }, \frac{ 1 }{ 1000 }\]

OpenStudy (anonymous):

@euler271 Thank you for explaining it.! I now understand it